# for Enhancing Visual Text Rendering

### v2

@@ Abstract
Visual Text Rendering (VTR) remains a critical challenge in text‑to‑image generation, where even advanced models frequently produce text with structural anomalies such as distortion, blurriness, and misalignment. However, we find that leading MLLMs and specialist OCR models largely fail to perceive these structural anomalies, creating a critical bottleneck for both VTR evaluation and RL‑based optimization. As a result, even state‑of‑the‑art generators (e.g., SeedDream4.0, Qwen‑Image) still struggle to render structurally faithful text. To address this, we propose TextPecker, a plug-and-play structural anomaly perceptive RL strategy that mitigates noisy reward signals and works with any text-to-image generator. To enable this capability, we construct a recognition dataset with character‑level structural‑anomaly annotations and develop a stroke‑editing synthesis engine to expand structural‑error coverage. Experiments show that TextPecker consistently improves diverse text‑to‑image models; even on the well‑optimized Qwen‑Image, it significantly yields average gains of 4% in structural fidelity and 8.7% in semantic alignment for Chinese text rendering, establishing a new state-of-the-art in high-fidelity VTR. Our work fills a gap in VTR optimization, providing a foundational step towards reliable and structural faithful visual text generation.
@@ 1 Introduction
We identify a critical bottleneck shared by both VTR evaluation and reinforcement learning process: a lack of fine-grained structural anomaly perception in rendered text. As illustrated in Fig. 1, OCR models [15, 53] and MLLMs [1, 40] are inherently ill-suited for this task. Their failures manifest in two primary ways: (1) Misinterpretation: They over-rely on linguistic priors to “correct” or hallucinate semantic content from structurally flawed text, thereby ignoring subtle glyph-level defects (e.g., stroke deletions, misalignments, or spurious attachments). (2) Invisibility: They often fail to detect or simply dismiss low-confidence text regions, such as those with significant blurriness or distortion, treating them as non-existent. Therefore, evaluators yield unreliable text‑accuracy estimates and fail to assess structural quality, and reward signals become misleading. As a direct result, even state-of-the-art generators (e.g., Qwen-Image [55], SeedDream4.0 [5]) still struggle to render structurally faithful text. Our quantitative analysis in Tab. 2 further substantiates this issue.
@@ 3.2.2 Structural Perceptive Data Construction
Step 1: Text-rich Image Generation. The initial step involves constructing a large-scale dataset of text-rich visual images, covering diverse structural error types. Specifically, for English text generation, we draw prompts from TextAtlas5M [49] and Lex-10k [66], and leverage multiple English-capable generative models (Anytext, Stable Diffusion v1-5 [43], Stable Diffusion 3.5 [16], Flux [28], SeedDream3.0 [19], Qwen-Image [55]) to generate such images. For Chinese text generation, we first sample a comprehensive text corpus from WanJuan1.0 [23], ensuring coverage of modern Chinese common characters We then use Qwen3-235B-A22B [60] to generate descriptions of various font styles, which are integrated with the corpus to form final prompts for models including Cogview4 [13], Kolors [27], SeedDream3.0 [19], and Qwen-Image [55].

### v3

@@ Abstract
Visual Text Rendering (VTR) remains a critical challenge in text‑to‑image generation, where even advanced models frequently produce text with structural anomalies such as distortion, blurriness, and misalignment. However, we find that leading MLLMs and specialist OCR models largely fail to perceive these structural anomalies, creating a critical bottleneck for both VTR evaluation and RL‑based optimization. As a result, even state‑of‑the‑art generators (e.g., Seedream4.0, Qwen‑Image) still struggle to render structurally faithful text. To address this, we propose TextPecker, a plug-and-play structural anomaly perceptive RL strategy that mitigates noisy reward signals and works with any text-to-image generator. To enable this capability, we construct a recognition dataset with character‑level structural‑anomaly annotations and develop a stroke‑editing synthesis engine to expand structural‑error coverage. Experiments show that TextPecker consistently improves diverse text‑to‑image models; even on the well‑optimized Qwen‑Image, it significantly yields average gains of 4% in structural fidelity and 8.7% in semantic alignment for Chinese text rendering, establishing a new state-of-the-art in high-fidelity VTR. Our work fills a gap in VTR optimization, providing a foundational step towards reliable and structural faithful visual text generation.
@@ 1 Introduction
We identify a critical bottleneck shared by both VTR evaluation and reinforcement learning process: a lack of fine-grained structural anomaly perception in rendered text. As illustrated in Fig. 1, OCR models [15, 53] and MLLMs [1, 40] are inherently ill-suited for this task. Their failures manifest in two primary ways: (1) Misinterpretation: They over-rely on linguistic priors to “correct” or hallucinate semantic content from structurally flawed text, thereby ignoring subtle glyph-level defects (e.g., stroke deletions, misalignments, or spurious attachments). (2) Invisibility: They often fail to detect or simply dismiss low-confidence text regions, such as those with significant blurriness or distortion, treating them as non-existent.
Therefore, evaluators yield unreliable text‑accuracy estimates and fail to assess structural quality, and reward signals become misleading. As a direct result, even state-of-the-art generators (e.g., Qwen-Image [55], Seedream4.0 [5]) still struggle to render structurally faithful text. Our quantitative analysis in Tab. 2 further substantiates this issue.
@@ 3.2.2 Structural Perceptive Data Construction
Step 1: Text-rich Image Generation. The initial step involves constructing a large-scale dataset of text-rich visual images, covering diverse structural error types. Specifically, for English text generation, we draw prompts from TextAtlas5M [49] and Lex-10k [66], and leverage multiple English-capable generative models (Anytext, Stable Diffusion v1-5 [43], Stable Diffusion 3.5 [16], Flux [28], Seedream3.0 [19], Qwen-Image [55]) to generate such images. For Chinese text generation, we first sample a comprehensive text corpus from WanJuan1.0 [23], ensuring coverage of modern Chinese common characters We then use Qwen3-235B-A22B [60] to generate descriptions of various font styles, which are integrated with the corpus to form final prompts for models including Cogview4 [13], Kolors [27], Seedream3.0 [19], and Qwen-Image [55].